At 100 o C, the gaseous reaction A ® 2B + C is found to be first order. On starting pure A, it is found that at the end of 10 minutes the total pressure of system is 176.0 mm of Hg. After a long time the pressure = 270.0 mm Hg. The partial pressure of A at end of 10 minutes is (in mm of Hg) (A) 43 (B) 47 (C) 94 (D) 86



At 100oC, the gaseous reaction A ® 2B + C is found to be first order. On starting pure A, it is found that at the end of 10 minutes the total pressure of system is 176.0 mm of Hg. After a long time the pressure = 270.0 mm Hg. The partial pressure of A at end of 10 minutes is (in mm of Hg)  


        (A)    43


        (B)    47


        (C)    94


(D)    86


Grade:11

2 Answers

Radhika Batra
10 years ago

For the reaction,

 

                                A              2B     +      C

 

        Initial pressure    po             0              0

 

        After 10 min       Po−x          2x             x

 

        After long time    −              2po            po

 

Since p  no. of moles and after long time only B and C are formed, A has completely disappeared \ pressure of 270 mm is for 3 moles of gaseous products.

 

or p due to one mole = 270/3 = 90 mm

 

Hence initial pressure = 90 mm

 

\ pressure at 10 min = 176 = 90−x + 2x + x

 

or    x=(176-90)/2=43     mm

 

or Pressure due to A to 10 minutes = 90 – 43

 

                                                = 47 mm

naresh masaboina
10 years ago

b